OpenStudy (anonymous):

If (x + yi) + (2(x + yi) + 3i) = 9, what is x + yi?

OpenStudy (ikram002p):

let x+iy = Z z+(2Z+3i)=9 3Z+3i=9 Z+i=9 x+iy+i=9 x+(y+1)i=9 x=9 y+1=0---> y=-1

OpenStudy (chosenmatt):

= 3 - i

OpenStudy (ikram002p):

what is that @chosenmatt ?

OpenStudy (anonymous):

i dont undersand

OpenStudy (anonymous):

@ikram002p are you sure you are not hungry this time?

OpenStudy (ikram002p):

eh why :O its correct @ganeshie8 , right ?

OpenStudy (anonymous):

Because you have eaten \(3\) there.. :P

OpenStudy (ikram002p):

3 what why where how who :O

OpenStudy (anonymous):

\(3Z + 3i = 9\) \(Z+i= 9\) ???

OpenStudy (ikram002p):

oh my ! its like the worst mistake xD

OpenStudy (ikram002p):

its ok , then x=3 not 9 :P

OpenStudy (anonymous):

What : \(3\) Why : you have eaten, so you know the best. Where : on very first comment on this post. How : you only know. :P Who : YOU.. :P

ganeshie8 (ganeshie8):

thats one of the main reasons why i let the OP do most of the work ;p

OpenStudy (ikram002p):

haha lol xD u never mad such a mistake xD

OpenStudy (ikram002p):

made*

OpenStudy (anonymous):

OP?

OpenStudy (ikram002p):

and by cheking there is no need in x, y sub only z+i=3 z=3-i x+yi=3-i